# Adopt a PET! An Exploration of PETs, Policy, and Practicalities for Industry in Canada

### v1

@@ 4. Methodology
Our study is targeted at the needs, norms, and expectations of industry actors in Canada in regard to privacy regulation and privacy technologies. As the needs and expectations of industry actors in Canada regarding privacy technology adoption and the influence of regulation are not well known, we wanted to avoid prematurely narrowing the scope of permitted responses to something that would fit within a quantitative survey style. We also faced challenges with time availability and anonymity were we to do a focus group or interview style study. Thus, we chose to do an open-ended survey style to (i) ease access and effort for our intended participant pool while still (ii) affording freedom to our participants concerning how much or how little they shared with us. The final study consists of an exploratory online survey with six questions where our responses were collected between May 20 and January 2025.
@@
We investigate how the industry of Canada complies with existing policies, such as PIPEDA (pipeda2000, ), as well as how they are impacted or influenced by emergent privacy regulations such as C-27 (and the AIDA act) (billc27, ). Our survey questions are designed with a focus on (i) understanding the existing data handling practices and compliance procedures of companies in Canada, (ii) exploring the impact of regulatory frameworks on privacy practices and technological advancements, and (iii) identifying gaps and challenges in current privacy policies and technologies.
We focus on Canada as it has had a version of a national privacy law since the 1980s when the Privacy Act (theprivacyact, ) was first formulated, and has since seen revisions with PIPEDA in 2000, and a recently proposed update to PIPEDA that was a part of Bill C-27. Based on the privacy law evolution that exists in Canada, we hypothesize that Canada’s industries have some basic ways of handling sensitive customer or client data as part of their normal processes. We further hypothesized that the emergence of new regulations may result in companies having concerns that their current, previously compliant, data handling practices may be insufficient and are actively seeking resources and strategies that could address their concerns.
@@
We primarily recruited our participants through a non-university mailing list (a not-for-profit institute in the researcher’s country) as well as through the researcher’s own networks of industry folks. Two participants were recruited through the sharing of a QR code at an event about new regulations in Canada. The rest of the potential participants received an email, either from the researchers or from the non-profit institute. The email described the research and provided a link to our short survey. Following the link would bring potential participants to our consent information, and then, if they so choose, they could proceed to answer our survey questions or exit the link. Participants were informed clearly that there was no expectation or requirement to participate and that they could quit the study at any time during the survey. To ensure there was no power differential between the researchers and those recruited, the survey was completely anonymous such that we are unable to determine who did or did not complete the study.
@@
While we did not collect personal demographic data, we only recruited those who did satisfy our requirements. Participants were to be between the ages of 18 and 65, working in industry in Canada, and English speaking. We limit to adults working in Canada as the focus of our investigation is Canadian industry’s approach to compliance with privacy regulations. We restrict to English speaking as that is the language in which the study will be administered.
@@
Our initial recruitment via mailing lists resulted in responses from 16 participants. We then recruited an additional six participants using our personal industry networks. These additional six were selected from different industries and in different roles to test our proximity to reaching thematic saturation (glaser2017discovery, ), at which point after the six participants’ responses were included, we found that saturation was reached and we stopped recruitment at our total of $N=22$ participants. We assigned each participant in our final set a random identifier between one and 60, and referred to them throughout as P#, where # is their identifier.
@@
We analyzed participant responses using an inductive approach to allow themes to emerge (saldana2021coding, ). Responses were organized first by survey questions. Then, the team of researchers analyzed all participant’s responses to that question and collaboratively formed clusters, which became our sub-themes. Then, we analyzed these sub-themes together with the survey question themes viewed collectively with respect to the research questions that the survey questions were targeted at. This second phase of analysis produced the final themes that are stated in our results and presented as an overview in Table 1.
@@
Our study focuses on industry within Canada and thus has a Canada-based population, which correspondingly, is a WEIRD population (schulz2018origins, ). While this is a limitation of our study, it is a deliberate one to allow us to focus on regulations local to a region. Our responses may have been impacted by the timing of our data collection as most of our responses came in while Bill C-27, which was a privacy law, was under consideration by the Canadian government. We cannot predict how the potential changes that would have resulted had C-27 passed could have impacted our participants or how it may have made regulations more prevalent in participants’ minds. Correspondingly, we cannot know how it could have impacted the participants’ perceptions on what responses would be more socially desirable, a factor that persists in human-response based empirical studies (redmiles2018asking, ).
@@
Theme: Process descriptions include a reliance on external organizations rather than specific PETS for privacy compliance
@@ 9. Discussion
Fischer et al. (Fischer2024Challenges, ) provide a visualization of the cryptography ecosystem and the path of bringing cryptography from paper to product. While the technologies are intertwined, there is a significant distinction between the journey of cryptography and privacy. In the former, end users and organizations are mostly the consumers of the product, where the final arrows point to. For privacy however, all elements of the socio-economic system — from regulations to client demand, consultants, competitors, best practices, market position, and reputation — play an active role in various phases of the decision-making journey. We associate this difference to the nature of privacy as a social concept, and the public awareness of digital technology and its potential threats and misuses in the current era. Hence, we observe these actions unfolding simultaneously: regulators striving to translate public demand into legislation, industry working to meet market privacy needs and ensuring compliance, and researchers developing privacy-preserving technologies that align with real-world concerns by supporting regulation and industry adoption. This interdependence can be frustrating, as each group relies on others whose work is still evolving. However, with effective communication, these efforts can align and reinforce one another.
Compliance verification is an example of a pain point that is notably challenging at this time. To elaborate on the nature of the compliance issue, we focus on the example of anonymity and de-identification as a fundamental pillar of privacy regulations. On one hand, de-identification and anonymity is one of the main privacy topics that were brought up repeatedly in our survey responses. Many mentioned it to be one of the widely deployed privacy enhancing technologies/methods in their industry. However, as we mentioned in Section 1, Canadian regulators are highly concerned about industry self regulating (ElEmam2024Perspectives, ) on the topic, since the methods industry uses for anonymity are not clear in design or implementation, and are not unified across industry. On the other hand, there are numerous research papers on how various PETs can assist with anonymity and de-identification or evaluate it. This presents an opportunity for collaboration among the three parties. Research institutes, in partnership with industry leaders, can respond to Canadian regulators’ call for issuing codes of practice and certifications by non-regulatory bodies. This can be achieved by leveraging PETs to provide a unified, systematic, and quantifiable evaluation of anonymization, such as through the use of privacy attacks such as membership inference (stadler2022synthetic, ; Giomi2023AUnified, ).
@@ 9.1. Key findings by stakeholder
It is advantageous that the usability of PETs is already receiving a lot of attention in the research community. However, as our participants pointed out, it is also crucial to consider PETs compatibility and integration with the existing platforms in industry, along with their associated cost as discussed (recall from Section 6). Furthermore, since regulations can accelerate the drive towards PETs adoption within industry, it would be highly beneficial to develop PETs that can be verified, audited, and demonstrably compliant with relevant regulations (recall from Section 8). The issuance of certificates by research institutes would be a collaborative effort that could help with the verification process, particularly if it aligned with regulatory requirements, which could then further support PETs adoption. The development of PETs that assist in evaluating privacy requirements will also be greatly valuable to industry actors as it will aid in facilitating regulatory enforcement with additional transparency and testability.
Understanding market dynamics in PETs adoption can help researchers navigate the challenges of bringing their work into industry. Our survey suggests that the choice of which PETs to adopt may be influenced by recommendations from existing service providers. Therefore, a viable approach to enhance acceptance is integrating PETs with established solutions. Another market factor that can affect the adoption process is company size. Smaller companies may implement PETs more quickly due to management involvement earlier in the process, but they may struggle to convince clients to adopt solutions that lack widespread use among larger companies. Additionally, we identify the personnel departments responsible for assessing the need for PETs and exploring mitigation strategies, which are not limited to personnel on privacy focused teams. Targeted education and effective communication with these departments is advised for any efforts aimed at significantly improving adoption rates.
@@
Our study suggests that industry professionals often seek clearer, more accessible communication regarding the regulations that apply to them, along with practical advice on how to achieve compliance. We note that according to a recent Canadian business survey result (phoenixspi2024survey, ), only one out of four of Canadian businesses have used the information and tools provided by OPC (Office of the Privacy Commissioner of Canada) despite being aware of them. In our survey, respondents (recall Section 5) reported government resources to be challenging to use and that they refer to privacy newsletters and other more easily understood sources when trying to decipher new requirements. Informed by our findings, we suggest that it would be beneficial to offer guidance that combines the authority of official sources with the actionable insights typically found in informal channels such as blogs, newsletters, or technical tutorials. Our survey identifies the key resources that Canadian industries rely on for regulatory compliance; from domain experts to consulting firms. Privacy regulators may find it more effective to address communication challenges by engaging with these intermediary parties, to ensure that industry receives clearer and more actionable compliance guidance. This approach could not only improve clarity but also make compliance more attainable for smaller companies that lack dedicated teams for separate tasks or face challenges in securing external counsel, which may lead them to adopt a follower strategy based on larger companies’ practices.
@@
While no outcome is more rewarding for researchers than seeing their designed PETs deployed in practice, this deployment remains a challenge. Our findings highlight key privacy focus areas in industry that we suggest will benefit from existing research dedicated to PETs design and development in these areas. While industry participants in our study employ well-established products, emerging PETs developed by researchers can offer valuable contributions; both in terms of privacy-preserving and privacy evaluation solutions. For example, in anonymity and de-identification context, current commercial data synthesis applications (Patki2016synthetic, ; gretel, ) either follow data sanitizing practices such as removing identifiers, or have not yet reached a level of maturity that allows them to scale effectively with the complexities and volume of industry data. However, there are several research papers advancing against the various practical problems of data synthesis (kollovieh2023selfguiding, ; pang2024clavaddpmm, ; zhang2023mixed, ).
@@ 9.2. Future Research Directions
In the context of anonymity, differential privacy has been successfully implemented by the U.S. Census Bureau and several large companies. Similarly, synthetic data generation has gained popularity in finance and health sectors. JP Morgan’s recent 2024 report (JPMorgan2024synthetic, ) advocates for the use of synthetic data in finance applications, highlighting several benefits, including enhanced privacy. Additionally, a national AI institute in Canada hosted a synthetic data generation bootcamp for numerous Canadian businesses at their request. However, these technologies were not explicitly mentioned by our participants. It remains unclear to us whether these technologies are not being utilized by the businesses in our survey, or if the participants themselves are unaware of them. A future research could help identify where/whether the state-of-the-art PETs are being deployed and clarify their adoption across industries. This would provide valuable insights for both researchers and industry practitioners.
@@
Future research on the effect of AI on privacy critical areas is crucial as AI technologies continue to evolve and integrate into various sectors. This effect is twofold. On one hand, it can complicate privacy compliance. For example, the common strategy of “removing identifiers” becomes even less effective in the age of AI, as its advanced pattern recognition and memorization capabilities (10.1145/3357713.3384290, ) enables re-identification more easily than ever before. Similarly, consent tracking or data disposal could become more challenging when the data is used in model training. On the other hand, AI-driven solutions may assist with compliance; synthetic data produced by generative models that can replace the original sensitive data is one such example.

### v2

@@ 4. Methodology
Our study is targeted at the needs, norms, and expectations of industry actors in Canada in regard to privacy regulation and privacy technologies. As the needs and expectations of industry actors in Canada regarding privacy technology adoption and the influence of regulation are not well known, we wanted to avoid prematurely narrowing the scope of permitted responses to something that would fit within a quantitative survey style. We also faced challenges with time availability and anonymity were we to do a focus group or interview style study. These challenges were communicated to us during our recruitment effort as being based in industry’s information sharing restrictions for their employees. Thus, we chose to do an open-ended survey style to (i) ease access and effort for our intended participant pool while still (ii) affording freedom to our participants concerning how much or how little they shared with us. In the end, our survey facilitated a non-presumptuous understanding of industry’s perspective on privacy, and allowed us to point out areas for future research. The final study consists of an exploratory online survey with six questions where our responses were collected between May 2024 and January 2025.
@@
We investigate how the industry of Canada approaches the adoption of novel privacy technologies and what impacts this adoption with consideration as to the potential relationship of policies such as PIPEDA (of Canada, 2000) and Bill C-27 (and the AIDA act) (of Canada, 2022). Our survey questions are designed with a focus on (i) understanding the existing data handling practices and procedures of companies in Canada, (ii) exploring the impact of regulatory frameworks on privacy practices and technological advancements, and (iii) identifying gaps and challenges in current privacy policies and technologies.
We focus on Canada as it has had a version of a national privacy law since the 1980s when the Privacy Act (of Canada, 1985) was first formulated, and has since seen revisions with PIPEDA in 2000, and a recently proposed update to PIPEDA that was a part of Bill C-27. Based on the privacy law evolution that exists in Canada, we hypothesize that Canada’s industries have some basic ways of handling sensitive customer or client data as part of their normal processes. We further hypothesized that the emergence of new regulations may result in companies having concerns that their current, previously compliant, data handling practices may be insufficient and are actively seeking resources and strategies that could address their concerns.
@@
In developing our questions, we ensured that our design investigates what technologies the industry perceives as privacy-preserving, even if they don’t match what research experts in privacy would classify as PETs. We have made an intentional effort in our survey not to predefine these notions, thereby reducing bias. Finally, before releasing our study we ran a small pilot on “similar” relevant folks prior to releasing the survey as part of our development process. No issues emerged with the study during this pilot.
@@
We primarily recruited our participants through a non-university mailing list (a not-for-profit institute in the researcher’s country) as well as through the researcher’s own networks of industry folks. This email list includes 170 contacts of which 143 engaged with the email in some fashion. Of those that engaged with the email approximately 55% have under 200 employees and the remainder have over 200 employees. Of those contacted, the majority are in the information technology sector followed by the health sector and then the financial sector. Two participants were recruited through the sharing of a QR code at an event about new regulations in Canada where the majority of participants were from companies in the technology or financial sector. The rest of the potential participants, recruited for saturation testing, received an email from the researchers. We contacted seven potential participants for our saturation check of which five were non-senior in their careers while two were senior decision makers. Among the seven contacted, three were from the financial sector and three from the technology sector. Finally, five of the total potential participants contacted in the check stage have privacy technology experience and in terms of customer size three were from organizations with under 200 employees while four were from organizations with over 200 employees.
@@
The email, whether from the researchers or sent via the non profit organization, described the research and provided a link to our short survey. Following the link would bring potential participants to our consent information, and then, if they so choose, they could proceed to answer our survey questions or exit the link. Participants were informed clearly that there was no expectation or requirement to participate and that they could quit the study at any time during the survey. To ensure there was no power differential between the researchers and those recruited, the survey was completely anonymous such that we are unable to determine who did or did not complete the study.
@@
While we did not collect personal demographic data, we only recruited those who did satisfy our requirements. Participants were to be between the ages of 18 and 65, working in industry in Canada, and English speaking. We limit participants to adults working in Canada as the focus of our investigation is the Canadian industry’s approach to compliance with privacy regulations. We restrict to English speaking as that is the language in which the study will be administered.
@@
Our initial recruitment via mailing lists resulted in responses from 16 participants. We then recruited an additional six participants using our personal industry networks. These additional six were selected from different industries and in different roles to test our proximity to reaching thematic saturation (Glaser and Strauss, 2017), at which point after the six participants’ responses were included, we found that saturation was reached and we stopped recruitment at our total of $N=22$ participants. We assigned each participant in our final set a random identifier between one and 60, and referred to them throughout as P#, where # is their identifier.
We note that being a privacy expert is not a requirement for our participants as we are trying to understand industry’s perceptions and beliefs about PETs and not to assess their expertise. Rather than evaluating whether or not “industry” has the correct views and whether they are “qualified” to speak on these things we focused on hearing openly what industry members actually think about and respond. Quality of responses was very rich with participants providing several sentences to paragraphs per question as their responses.
Research team.
Our team includes industry experience as a cryptography consultant as well as expertise and experience in qualitative coding and security and privacy technical expertise. Collectively our team holds expertise relevant to creating and assessing PETS, industry and PETs, and qualitative research.
@@
We analyzed participant responses using an inductive approach to allow themes to emerge (Saldaña, 2021). Responses were organized first by survey questions. Then, the team of researchers analyzed all participant’s responses to that question and collaboratively formed clusters, which became our sub-themes. Then, we analyzed these sub-themes together with the survey question themes viewed collectively with respect to the research questions that the survey questions were targeted at. This second phase of analysis produced the final themes that are stated in our results and presented as an overview in Table 1.
@@
Our study focuses on industry within Canada and thus has a Canada-based population, which correspondingly, is a WEIRD population (Schulz et al., 2018). While this is a limitation of our study, it is a deliberate one, to allow us to focus on an understudied region with established and emerging privacy norms and regulations. Our responses may have been impacted by the timing of our data collection as most of our responses came in while Bill C-27, which was a privacy law, was under consideration by the Canadian government. We cannot predict how the potential changes that would have resulted had C-27 passed could have impacted our participants or how it may have made regulations more prevalent in participants’ minds. Correspondingly, we cannot know how it could have impacted the participants’ perceptions on what responses would be more socially desirable, a factor that persists in human-response based empirical studies (Redmiles et al., 2018).
@@
Theme: Process desc.s include a reliance on external organizations rather than specific PETS for privacy compliance
@@ 6. Results on factors for PETs adoption
[…]
[Flows from step i to iv with sub-steps described]Shows decision-making process for PETs adoption as captured from the respondents. The process flows from top to bottom for the large rectangular steps and proceeds through the different components within each larger overall step.
@@ 9. Discussion
Fischer et al. (Fischer et al., 2024) provide a visualization of the cryptography ecosystem and the path of bringing cryptography from paper to product. While the technologies are intertwined, there is a significant distinction between the journey of cryptography and privacy. In the former, end users and organizations are mostly the consumers of the product, where the final arrows point to. For privacy however, all elements of the socio-economic system — from regulations to client demand, consultants, competitors, best practices, market position, and reputation — play an active role in various phases of the decision-making journey. We associate this difference to the nature of privacy as a social concept, the public awareness of digital technology, and its potential threats and misuses in the current era. Hence, we observe these actions unfolding simultaneously: regulators striving to translate public demand into legislation, industry working to meet market privacy needs and ensuring compliance, and researchers developing privacy-preserving technologies that align with real-world concerns by supporting regulation and industry adoption. This interdependence can be frustrating, as each group relies on others whose work is still evolving. However, with effective communication, these efforts can align and reinforce one another.
Compliance verification is an example of a pain point that is notably challenging at this time. To elaborate on the nature of the compliance issue, we focus on the example of anonymity and de-identification as a fundamental pillar of privacy regulations. On one hand, de-identification and anonymity is one of the main privacy topics that were brought up repeatedly in our survey responses. Many mentioned it to be one of the widely deployed privacy enhancing technologies/methods in their industry. However, as we mentioned in Section 1, Canadian regulators are highly concerned about industry self regulating (El Emam et al., 2024) on the topic, since the methods industry uses for anonymity are not clear in design or implementation, and are not unified across industry. On the other hand, there are numerous research papers on how various PETs can assist with anonymity and de-identification or evaluate it. This presents an opportunity for collaboration among the three parties. Research institutes, in partnership with industry leaders, can respond to Canadian regulators’ call for issuing codes of practice and certifications by non-regulatory bodies. This can be achieved by leveraging PETs to provide a unified, systematic, and quantifiable evaluation of anonymization, such as through the use of privacy attacks like membership inference (Stadler et al., 2022; Giomi et al., 2022).
@@ 9.1. Key findings by stakeholder
It is advantageous that the usability of PETs is already receiving a lot of attention in the research community. However, as our participants pointed out, it is also crucial to consider PETs compatibility and integration with the existing platforms in industry, along with their associated cost as discussed (recall from Section 6). Since regulations can accelerate the drive towards PETs adoption within industry, it would be highly beneficial to develop PETs that can be verified, audited, and demonstrably compliant with relevant regulations (recall from Section 8). The issuance of certificates by research institutes would be a collaborative effort that could help with the verification process, particularly if it aligned with regulatory requirements, which could then further support PETs adoption. The development of PETs that assist in evaluating privacy requirements will also be greatly valuable to industry actors as it will aid in facilitating regulatory enforcement with additional transparency and testability.
Understanding market dynamics in PETs adoption can help researchers navigate the challenges of bringing their work into industry. Our survey suggests that the choice of which PETs to adopt may be influenced by recommendations from existing service providers. While this result parallels many attributes identified in prior work on the value of networks as professional connections for knowledge or expertise sharing (Ehrlich et al., 2007), competitors and best practices were unique factors that were highlighted as ways new techniques come to industry’s awareness and subsequent adoption. This means that not only is it the case that social influence impacts privacy behaviors of those interacting with social networks (Mendel and Toch, 2017), but also that the social hierarchies of who organizations view as peers, newcomers to the field, and leaders in the field all can impact organizations in their inclination to adopt privacy technologies. Similarly, smaller companies may implement PETs more quickly due to management involvement earlier in the process, but they may struggle to convince clients to adopt solutions that lack widespread use among larger companies. Additionally, we identify the personnel departments responsible for assessing the need for PETs and exploring mitigation strategies, which are not limited to personnel on privacy focused teams. This suggests that an effective way of propagating PETs in industry is illustrating who among their peers uses the same or similar systems. Such a peer based grouping seems to be the closest approximation to communities of practice for our participants (Wenger, 1999). Therefore, a viable approach to enhance acceptance is integrating PETs with established solutions. As participants identify blogs and digital newsletters as a source of information, an example of computer-mediated communication that could be harnessed by the technical research community more effectively (Preece, 2000), such platforms may be effective if used to communicate about usages and acceptance rates by peers.
@@
Our study suggests that industry professionals often seek clearer, more accessible communication regarding the regulations that apply to them, along with practical advice on how to achieve compliance. We note that according to a recent Canadian business survey result (Phoenix Strategic Perspectives Inc., 2024), only one out of four of Canadian businesses have used the information and tools provided by OPC (Office of the Privacy Commissioner of Canada) despite being aware of them. In our survey, respondents (recall Section 5) reported government resources to be challenging to use and that they refer to privacy newsletters and other more easily understood sources when trying to decipher new requirements. Informed by our findings, we suggest that it would be beneficial to offer guidance that combines the authority of official sources with the actionable insights typically found in informal channels such as blogs, newsletters, or technical tutorials. Our survey identifies the key resources that Canadian industries rely on for regulatory compliance; from domain experts to consulting firms. Privacy regulators may find it more effective to address communication challenges by engaging with these intermediary parties, to ensure that industry receives clearer and more actionable compliance guidance. This approach could not only improve clarity but also make compliance more attainable for smaller companies that lack dedicated teams for separate tasks or face challenges in securing external counsel, which may lead them to adopt a follower strategy based on larger companies’ practices.
@@
While no outcome is more rewarding for researchers than seeing their designed PETs deployed in practice, this deployment remains a challenge. Our findings highlight key privacy focus areas in industry that we suggest will benefit from existing research dedicated to PETs design and development in these areas. While industry participants in our study employ well-established products, emerging PETs developed by researchers can offer valuable contributions; both in terms of privacy-preserving and privacy evaluation solutions. For example, in anonymity and de-identification context, current commercial data synthesis applications (Patki et al., 2016; AI, nd) either follow data sanitizing practices such as removing identifiers, or have not yet reached a level of maturity that allows them to scale effectively with the complexities and volume of industry data. However, there are several research papers advancing against the various practical problems of data synthesis (Kollovieh et al., 2023; Pang et al., 2024; Zhang et al., 2024).
@@ 9.2. Future Research Directions
In the context of anonymity, differential privacy has been successfully implemented by the U.S. Census Bureau and several large companies. Similarly, synthetic data generation has gained popularity in finance and health sectors. JP Morgan’s recent 2024 report (Potluru et al., 2024) advocates for the use of synthetic data in finance applications, highlighting several benefits, including enhanced privacy. Additionally, a national AI institute in Canada hosted a synthetic data generation bootcamp for numerous Canadian businesses at their request. However, these technologies were not explicitly mentioned by our participants. It remains unclear to us whether these technologies are not being utilized by the businesses in our survey, or if the participants themselves are unaware of them. A future research could help identify where/whether the state-of-the-art PETs are being deployed and clarify their adoption across industries. This would provide valuable insights for both researchers and industry practitioners.
@@
Future research on the effect of AI on privacy critical areas is crucial as AI technologies continue to evolve and integrate into various sectors. This effect is twofold. On one hand, it can complicate privacy compliance. For example, the common strategy of “removing identifiers” becomes even less effective in the age of AI, as its advanced pattern recognition and memorization capabilities (Feldman, 2020) enables re-identification more easily than ever before. Similarly, consent tracking or data disposal could become more challenging when the data is used in model training. On the other hand, AI-driven solutions may assist with compliance; synthetic data produced by generative models that can replace the original sensitive data is one such example.
@@ 10. Conclusion
Through this study, we identify the breadth of approaches employed by organizations considering PETs and the challenges they face. We further identify a gap between how companies think of privacy technologies and how researchers think of privacy technologies that can contribute to low adoption of the increasingly sophisticated privacy technologies produced by researchers, such as applications of differential privacy, multiparty computation, and trusted execution environments.
@@
Acknowledgments
The work of Xi He was supported by NSERC through a Discovery Grant, an alliance grant, and the Canada CIFAR AI Chairs program. The work of Bailey Kacsmar was supported by NSERC through a Discovery Grant (RGPIN-2024-04996).